- Press Release
- Sep 26, 2022
NOAA GOES SATOPS Morning Report 1 November 2000
SATOPS MORNING REPORT November 1, 2000
No change in the status of the POES spacecraft.
NOAA-16 operations nominal.
Due to reports that AMSU products have been significantly degraded since
the NOAA-15 solar array movement (and subsequent cooling down of the AMSU
instruments) from -45 degrees to +53 degrees on October 30 to facilitate
AVHRR troubleshooting. As a result, the array will be returned to
the original offset of -45 degrees at the earliest opportunity today, tentatively
scheduled for Rev 12843/Fairbanks, which at 16:40z today. We plan
to complete the remainder of the AVHRR tests scheduled for this week at
the -45 degree offset.
NOAA-16 rev 567 / F at 1509z on Tuesday, October 31: the solar
array offset was changed from -35 degrees to -40 degrees.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 1942z on Tuesday, October 31, a GOES-10 frame break was observed
and attributed to loss of signal sync.
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.
GOES operations were nominal over the past 24 hours.
Questions or comments to email@example.com